OpenStudy (anonymous):

MINIMUM AND MAXIMUM remind me how to find the min and max?!? y=-5cos(x+3)-1

OpenStudy (anonymous):

where do you think the middle line is?

OpenStudy (anonymous):

the maximum of cosine is 1 the maximum of 5 times cosine is therefore 5 and so the minimum of - 5 times cosine is -1 therefore, the minimum of \(-5\cos(x)-1\) is ...

OpenStudy (anonymous):

oops, i meant to say " the minimum of -5 times cosine is -5"

OpenStudy (anonymous):

im looking at a graph of it and that's a very good explanation im just very lost it looks like the maximum is 4

OpenStudy (anonymous):

hope it is clear that the \(x+1\) in the input has no effect on the max or min

OpenStudy (anonymous):

yes, the max is 4 lets walk through it

OpenStudy (anonymous):

the minimum value of cosine is \(-1\) the minimum value of 5 times cosine is therefore \(-5\) and so the maximum value of \(-5\) times cosine is 5

OpenStudy (anonymous):

it looks like the minimum vale is -6 though

OpenStudy (anonymous):

since you are subtracting 1 at the end, the maximum value of \(-5\cos(x+1)-1\) is \(4\)

OpenStudy (anonymous):

yes, \(-6\) same reasoning as above

OpenStudy (anonymous):

oh okay that makes sense

OpenStudy (anonymous):

the minimum value of \(-5\cos(x+1)\) is \(-5\) and since you subtract 1 at the end, then the minimum value of \(-5\cos(x+1)-1\) is \(-5-1=-6\)

OpenStudy (anonymous):

thank you thank you